OpenStudy (anonymous):

Elise and Miguel both collect baseball cards. Miguel has 8 more than 4 times as many cards as Elise. Together they have 993 cards. How many cards does each person have?

OpenStudy (anonymous):

how many cards does each person have?

OpenStudy (anonymous):

Please help :(

OpenStudy (judygreeneyes):

OK, here it comes

OpenStudy (anonymous):

Thank you! I think I know the answer, but I'm not sure…

OpenStudy (anonymous):

if it's correct

OpenStudy (judygreeneyes):

Elise has X cards. Miguel has 4 times as many, plus 8, so he has 4X+8 cards. Together, they have X + 4X+8 = 993 cards. Now we solve for X. Combine terms and move the numbers over to the right side: 5X +8 = 993 5X = 985 X = 197 So Elise has 197 cards. This means that Miguel has 4*197 + 8 = 796 cards. Double check: 197 + 796 = 993 Yay!!

OpenStudy (anonymous):

Thank you so much! Now I know I got it right on my assignment :)